OpenStudy (anonymous):

Can somebody help me with these questions please? 1. Non-square matrices have inverses. Answer: Explanation: 2. Multiplication of a matrix and its inverse is commutative. Answer: Explanation: 3. In a system of linear equations, if the determinant of the coefficient matrix is zero, the system has no solution. Answer: Explanation: 4. If you multiply two matrices and obtain the identity matrix, you can assume the matrices are inverses of each other. Answer: Explanation:

OpenStudy (dan815):

wait let me see

OpenStudy (dan815):

non square matrices cannot have inverse

OpenStudy (dan815):

because the identity matrix is a square matrix

OpenStudy (dan815):

i know that left inverse means there exists a right inverse but i dont htink they necessairly have to be the same

OpenStudy (dan815):

3)it means there is a solution its invertible so it can be solved

OpenStudy (dan815):

4) yea be cause AB=I then A=B^-1 or B=A^-1 as B=A^-1*I = A^1 and A*B*B^-1 = I*B^-1 so A=B^-1

OpenStudy (anonymous):

omg . thanks so much ! :)